what is the electric field from an electron 1.0cm away from it?


1
Expert's answer
2021-04-13T06:36:06-0400

Electric field of point charge "q" at distance "r" from it is "E = \frac{k q}{r^2}", where "k \approx 9 \cdot 10^9 \frac{N \cdot m^2}{C^2}"is the Coulomb's constant.

Hence, "E = \frac{k e}{r^2} = \frac{9 \cdot 10^9 \frac{N \cdot m^2}{C^2} \cdot 1.6 \cdot 10^{-19} C}{(1 \cdot 10^{-2})^2 m^2} \approx 14.4 \cdot 10^{-6} \frac{V}{m}".